OpenStudy (anonymous):

Sketch the curve x^4 + y^4 = u Where u is any positive constant. We are supposed to be able to do this pretty easily from looking at the equation, but I'm not sure how, can anybody explain?

OpenStudy (anonymous):

Initially I said Let a = x^2 and b = y^2 then sketched a circle on the axes of a and b with radius sqrt(u) but I couldn't project that onto the xy plane

OpenStudy (anonymous):

The u is just a kind of scaling so u might as well put it as 1 in order to get the shape. After that put x= 1 and -1, go from there.

OpenStudy (anonymous):

Do what from there?

OpenStudy (anonymous):

Well, if x=1, y= plus/minus 1 etc and plot.

OpenStudy (anonymous):

etc? I don't know what you're saying

OpenStudy (anonymous):

@estudier ?

OpenStudy (anonymous):

Do you not know how to plot values on a graph?

OpenStudy (anonymous):

Do you realize there is more than one way to interpolate between points?

OpenStudy (anonymous):

Yes, I do. Why do you think this requires any form of interpolation?

OpenStudy (anonymous):

Because a graph is not a set of a dots.

OpenStudy (anonymous):

You want a sketch, right. A sketch is not a graph.

OpenStudy (anonymous):

Sorry I should've said "Sketch the graph of the curve"

OpenStudy (anonymous):

The curve is symmetric and continuous, smooth even. Plot a few points and the sketch is simple (evidently it is not a rectangle)

OpenStudy (anonymous):

The original question was about identifying the axes of symmetry of the curve, y=0 and x=0 are obvious, but there's more I can't figure out

OpenStudy (anonymous):

(because I don't know what the function looks like)

OpenStudy (anonymous):

Wait a minute, I will get you a picture.....

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

How can I know that it is 'more rectangular' than a circle from looking at the equation?

OpenStudy (anonymous):

Please? I seriously don't understand this. @estudier

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

Plot some points x = 0 for example